what is the first term ,when the second term is 3/4 and the fourth term is 3 of geometric sequence
Answer by JoelSchwartz(130)

y=the value of the first term
x=the common ratio between each number
yx=3/4
y=.75/x
yx^3=3
.75x^2=3
3x^2=12
x^2=4
x=2
2y=3/4
8y=3
y=3/8
The value of the first term is 3/8